OpenStudy (anonymous):

What is the rate of change for f(x) = –2 cos 4x – 3 on the interval from x = 0 to x = pi/4

OpenStudy (anonymous):

Answer choices A. 2/pi B.pi/2 C.16/pi D. pi/16

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

okay so plug in values for both 0 and π/4

OpenStudy (anonymous):

and then use (f(0)-f(π/4))/(0-4)

OpenStudy (anonymous):

don't forget to medal... any questions just message me

OpenStudy (anonymous):

Hey sorry my computer had to be restarted. How do I give you a medal?

OpenStudy (anonymous):

you just click best answer, and it medals

OpenStudy (anonymous):

@cramos725 could you help me a bit more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what do u need help on?

OpenStudy (anonymous):

I am new to this type of problem entirely so could you show me the plugging in part?

OpenStudy (anonymous):

okay sure... so f(x) = –2 cos 4x – 3 when you plug in 0, you have -2 cos(4(0)) -3 which is equal to -2 cos(0) -3 = -2(1)-3 =-2-3 =-5

OpenStudy (anonymous):

oh ok I see what I did wrong

OpenStudy (anonymous):

when you plug in π/4, you get -2(cos4(π/4))-3 so you get -2 cos(π) -3 =-2(-1)-3 =-1

OpenStudy (anonymous):

glad i could help

OpenStudy (anonymous):

and the denominator should be (0-π/4)

OpenStudy (anonymous):

i made a mistake typing it in earlier

OpenStudy (anonymous):

you get -5-(-1) which is -4 all over 0-π/4 -4/(-π/4) which is 16/π

OpenStudy (anonymous):

oh i see. Thank you so much I really appreciate yout time.